Set up a proportion for the following situation.It takes Todd 8 minutes to eat 5 pies at a pie eating contest. How many pies can Todd eat in 15 minutes?
Answer by stanbon(75887)

Set up a proportion for the following situation.It takes Todd 8 minutes to eat 5 pies at a pie eating contest. How many pies can Todd eat in 15 minutes?
---
pies/min = pies/min
---
x/15 = 5/8
----
x = 15(5/8)
---
x = 75/8
----
x = 9.375 pies
===================
Cheers,
Stan H.
===================